Given the utility function, U = x2y2, with the budget constraint , M = P1 x + P2 y , where U is the utility, x is the quantity of commodity one and Y is the quantity of commodity two and M is income of the consumer .


a. find the utility maximizing quantities of both commodities

b. if the P1= 2 and P2 and M =50, determine the specific quantities of both commodities

c determine whether utility is maximized or minimized



1
Expert's answer
2021-04-26T09:40:45-0400

"Max U=X^2Y^2"

"St. P_1X+P_2Y=M"

"L=X^2Y^2-\lambda""(P_1X+P_2Y-M)"

"dL/dX=2XY^2-\lambda""P_1=0 ......i"

"dL/dY=2X^2Y-\lambda""P_2 =0......ii"

"dL/d\lambda""=P_1X+P_2Y-M=0....iii"

Rearranging equation i and ii and dividing

"2XY^2\/2X^2Y=P_1\/P_2"

"Y=P_1X\/P_2.......iv"

Hence; "X=P_2Y\/P_1.....v"

Replacing equation iv and v in equation iii then

"P_1(P_2Y\/P_1)+P_2Y=M"

"2P_2Y=M"

"Y=M\/2P_2" .......Utility maximizing quantity of Y

Hence; "X=M\/2P_1......." Utility maximizing quantity of X

b) Specific Maximizing quantities given that P1 =2 P2 and M=50

"Y=50\/2(50) =0.5"

"X=50\/2(2)=12.5"

c) It is maximized since the consumer is consuming certain amounts of each commodity and not one commodity hence achieving maximum utility.